OpenStudy (anonymous):

If the efficiency of a heat engine that does 1.00 × 10^3 J of work is 70 percent, then what is the heat input? A. 1.43 × 10 J B. -7.00 × 10^2 J C. 9.30 × 10^2 J D. 1.43 × 10^3 J **Not sure! thanks!!:)

OpenStudy (michele_laino):

by definition, the efficiency is given by the subsequent formula: \[\eta = \frac{L}{Q}\]

OpenStudy (anonymous):

okay! so 70/1.00*10^3 ?

OpenStudy (michele_laino):

no, it is: \[Q = \frac{L}{\eta } = \frac{{{{10}^3}}}{{0.7}}\]

OpenStudy (michele_laino):

\[\Large Q = \frac{L}{\eta } = \frac{{{{10}^3}}}{{0.7}}\]

OpenStudy (anonymous):

ohh oops!! okay so then we get 1428.57 ? so is that choice D?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

yay!! thank you:)

OpenStudy (michele_laino):

thanks!! :)